OpenStudy (anonymous):

Find a quadratic equation with the two given numbers as solutions: −6+3i5√ and −6−3i5√

OpenStudy (anonymous):

sorry that is -6 + 3i sqt 5 and -6 - 3i sqt 5

OpenStudy (anonymous):

I think he means −6+3i√5 and −6−3i√5

OpenStudy (anonymous):

\[x^2 + \text{(sum of roots)}x + \text{(product of roots)}\]

OpenStudy (anonymous):

just write it out like ishaan said...simple

OpenStudy (anonymous):

Would the answer be \[x ^{2} + 10x + 25?\]

OpenStudy (anonymous):

sum = -12 product = 36 - (- 9*5 )

OpenStudy (anonymous):

so it would be \[x^{2} - 12x + 45?\]

OpenStudy (anonymous):

Im confused of the product part. where did the 36 come from?

OpenStudy (anonymous):

6x6 = 36, -3i x -3i = 9 i^2 = -9, because i^2 = -1 and the 5 is from sqrt {5} x sqrt {5}

OpenStudy (anonymous):

sorry, that should be -6 x -6 = 36

OpenStudy (anonymous):

so the answer would then be x^2 - 12x + (36 + 45)?

OpenStudy (anonymous):

correctimundo Droid d00d

OpenStudy (anonymous):

COOOL ty

OpenStudy (anonymous):

Math is a scary thing, to _everyone_ When you reach the limit to your frontier of understanding, it is one of the most frightening things on the planet to cross past that frontier But few things compare to the swelling of joy one can obtain when he gains understanding of a concept and can produce good work and solve stuff. Kind of like what you're feeling right now :P